OpenStudy (anonymous):

Solve for x: 5x + 8 = -32

OpenStudy (anonymous):

5x + 8 = -32 5x = -32 -8 5x = -40/5 x = -8

OpenStudy (anonymous):

thank you!!!:)

OpenStudy (anonymous):

ur welcome!!!!